OpenStudy (anonymous):

Help!? In the study of electricity, the formula V=I*R is used to relate the voltage (V), current (I) and resistance (R) of circuits. What is the value of I when V=9 and R=0.2

OpenStudy (anonymous):

Answer choices 1.8 4.5 18.0 45.0

OpenStudy (trojanpoem):

V = IR I = V/R = 9 / 0.2 = 90 / 2 = 45

OpenStudy (anonymous):

thank you

OpenStudy (trojanpoem):

You are welcome. To solve it your self write everything down and put in in the formula